The angle subtended by the thickness of a ¼-inch thick pencil at 15 feet from the eye. Give your answer in radians and in degrees
1
Expert's answer
2019-10-15T06:22:42-0400

The value of the angle in radians is "\psi = \frac{0.25\frac{1}{12}}{15} = 1.4\cdot10^{-3}" radians

The value of the angle in degrees is "\beta = \psi\frac{180}{3.14} = 0.08" degrees